OpenStudy (anonymous):

Need help understanding this equation. 3(5z-1)-2(z+6)=10(z+1)

OpenStudy (anonymous):

Hint: Distribute the number outside the bracket, into the bracket.

OpenStudy (anonymous):

Okay thank you. I have simplified the equation down to 13z-15=10Z+10. I have forgotten how to combine the z's onto one side of the =

OpenStudy (anonymous):

If you subtract a number from one side, you must also subtract the number from the other side. Likewise with division and multiplication. Example: 2x+9=4x+10 2x-4x+9-9=4x-4x+10-9

OpenStudy (anonymous):

So for that example, you'll get... -2x=1 You want to get x, not -2x so you need to divide... -2x/-2=1/-2 x=1/-2 Do you get how it works now? :)

OpenStudy (anonymous):

Ah I see now. Thanks so much

OpenStudy (anonymous):

:) You're welcome!